OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the points (3,-4)

OpenStudy (compassionate):

Hello! My name is Stephen and I will be your guide. So, we're given the slope -3 and the two points (3, -4) The next step is to plug those into the equation: y - y1 = m(x - x1) Where y1 and x1 are your points and m is your slope. Plug them in: y - (-4) = -3(x - 3) Now, can you simplify the equation?